# A classification of Prüfer domains of integer-valued polynomials on algebras

### v1

@@ 1 Introduction
Integer-valued polynomials have a long history in commutative ring theory. One problem that attracted significant interest [4, 13, 5, 17] was to classify the domains $D$ such that $\textnormal{Int}(D)$ is a Prüfer domain. Necessary and sufficient conditions on $D$ for this to occur were determined in [18]. Since the late 2000s, attention has turned to rings of the form $\textnormal{Int}_{K}(A)$ and $\textnormal{Int}_{K}(S,A)$ [33, 19, 7, 9, 10, 22, 8, 11, 14, 20, 21]. Here, as in the traditional case where $A=D$, researchers have studied situations in which these rings are or are not Prüfer. For instance, when $D=\mathbb{Z}$ and $A$ is a ring of algebraic integers, it is known that $\textnormal{Int}_{\mathbb{Q}}(A)$ is Prüfer [19, Theorem 3.7]. By contrast, if $D$ is a Dedekind domain with finite residue fields and $A$ is the matrix algebra $A=M_{n}(D)$ with $n\geq 2$, then $\textnormal{Int}_{K}(A)$ is never integrally closed [28, Corollary 3.4], hence cannot be Prüfer.
@@
  • The domain $D$ contains a principal ideal domain $D_{0}$, the structure of which depends on the characteristic of $K$. If $\text{char}(K)=0$, then $D_{0}\cong\mathbb{Z}$. If $\text{char}(K)=p$, for some integral prime $p$, then $D_{0}\cong\mathbb{F}_{p}[t]$, the univariate polynomial ring over the finite field $\mathbb{F}_{p}$.
  • Let $\mathfrak{p}$ denote a nonzero prime element of $D_{0}$. Explicitly, if $D_{0}\cong\mathbb{Z}$, then $\mathfrak{p}$ is an integral prime; otherwise, $\mathfrak{p}$ is a polynomial irreducible over $\mathbb{F}_{p}$.
@@
  • For each $\mathfrak{p}$ and each $P$ over $\mathfrak{p}$, define
    $\quad e_{P,\mathfrak{p}}=v_{P}(\mathfrak{p})\quad\text{ and }\quad f_{P,\mathfrak{p}}=[(D/P):(D_{0}/\mathfrak{p}D_{0})].$
  • For each $\mathfrak{p}$, define
    $E_{\mathfrak{p}}:=\{e_{P,\mathfrak{p}}\mid P|\mathfrak{p}D_{0}\}\quad\text{ and }\quad F_{\mathfrak{p}}:=\{f_{P,\mathfrak{p}}\mid P|\mathfrak{p}D_{0}\}.$
  • We say that $D$ satisfies the double-boundedness condition if, for each $\mathfrak{p}$, both $E_{\mathfrak{p}}$ and $F_{\mathfrak{p}}$ are bounded.
@@ 3 Necessary conditions
In this section, we will establish necessary conditions for $\textnormal{Int}_{K}(A)$ to be Prüfer. Since one of our standard assumptions on $A$ is that $A\cap K=D$, we have $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}(D)$. Thus, if $\textnormal{Int}_{K}(A)$ is Prüfer, then so is $\textnormal{Int}(D)$. If the reader wishes, for the remainder of this section they may consider $D$ to be an ADDB-domain (i.e. $D$ is such that $\textnormal{Int}(D)$ is Prüfer), since this is the assumption we will make in our major theorems. However, we state our intermediary results using only the properties of the integrally closed domain $D$ that are needed to reach the desired conclusion.
@@
For any ring $R$, let $\mathscr{J}(R)$ be the Jacobson radical of $R$. We recall that $R$ is semiprimitive if $\mathscr{J}(R)=\{0\}$. If $D$ is semiprimitive (which occurs, for instance, when $D=\mathbb{Z}$), we can impose even stronger necessary conditions on $A$ in order for $\textnormal{Int}_{K}(A)$ to be Prüfer. To do this, we first show that in some cases the existence of a nonzero nilpotent element in a simple residue ring of $A$ is enough to force $\textnormal{Int}_{K}(A)$ to not be Prüfer.
@@
Let $D$ be a Prüfer domain and assume that $A$ contains a maximal ideal $\mathfrak{m}$ such that $A/\mathfrak{m}\cong M_{n}(F)$, where $n\geq 2$ and $F$ is a finite field. Then, $\textnormal{Int}_{K}(A)$ is not Prüfer.
@@
Let $P=D\cap\mathfrak{m}$. We first verify that $P$ is a prime ideal of $D$. Certainly, $P$ is a proper ideal of $D$ because $1\notin P$. Let $cd\in P$. Then, $cd\in\mathfrak{m}$, and since both $c$ and $d$ are central in $A$, we have $cAd\subseteq\mathfrak{m}$. The ring $A$ is noncommutative, but maximal ideals in noncommutative rings are still prime [16, p. 165], so the condition $cAd\subseteq\mathfrak{m}$ implies that either $c\in\mathfrak{m}$ or $d\in\mathfrak{m}$ [16, Prop. 10.2]. Thus, either $c$ or $d$ is in $P$.
Next, recall that $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}(D)$. If $P$ has infinite index in $D$, then by [2, Proposition I.3.4] we have $\textnormal{Int}(D)\subseteq D_{P}[X]$ and so $\textnormal{Int}_{K}(A)$ cannot be Prüfer. So, assume $D:P]<\infty$ and consider the $D_{P}$-algebra $A_{P}=(D\setminus P)^{-1}A$. Since $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}_{K}(A_{P})$, it suffices to show that $\textnormal{Int}_{K}(A_{P})$ is not Prüfer. Because $D$ is Prüfer, $D_{P}$ is a valuation domain. If $PD_{P}$ is not principal, then by [2, Proposition I.3.16] we have that $\textnormal{Int}(D_{P})=D_{P}[X]$, so the subring $\textnormal{Int}_{K}(A_{P})$ cannot be Prüfer. Suppose now that $PD_{P}$ is principal and let $\pi\in D_{P}$ be a generator of $PD_{P}$ Then, $A_{P}/\pi A_{P}$ has $M_{n}(F)$ as a residue ring and $D_{P}:\pi D_{P}]=[D:P]$ is finite. By Lemma 3.5, $\textnormal{Int}_{K}(A_{P})$ is not Prüfer, and we are done. ∎
@@
Assume that $\textnormal{Int}_{K}(A)$ is Prüfer. Then, $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}(D)$ because $A\cap K=D$, and so $\textnormal{Int}(D)$ is Prüfer. By [4, Prop. 6.3], $D$ must be an almost Dedekind domain with finite residue fields, hence $D$ is a Prüfer domain. Since $\textnormal{Int}_{K}(A)$ is Prüfer, the condition of Theorem 3.6 cannot hold. Therefore, every simple residue ring of AA must be a finite field. ∎
@@
In what follows, for any ring $R$, let $\textnormal{Max}(R)$ be the set of maximal ideals of $R$, called the maximal spectrum of $R$.
@@
Assume that $D$ is one-dimensional.
  1. (1)
    Let $\mathfrak{m}\in\textnormal{Max}(A)$. Then, $\mathfrak{m}\cap D\in\textnormal{Max}(D)$.
  2. (2)
    Let $M\in\textnormal{Max}(D)$. Then, there exists a maximal ideal of $A$ containing $MA$.
@@
(1) Let $M=\mathfrak{m}\cap D$. Working as in the proof of Theorem 3.6, we see that $M$ is a prime ideal of $D$. It remains to show that $M\neq\{0\}$. If this were the case, then $D$ would be isomorphic to a subring of $A/\mathfrak{m}$, because
$D=D/(\mathfrak{m}\cap D)\cong(D+\mathfrak{m})/\mathfrak{m}\subseteq A/\mathfrak{m}.$
However, $A/\mathfrak{m}$ is a simple ring and $D$ is central in $A$, so this would force $D$ to be a field, which is a contradiction. Thus, $M$ is a nonzero prime ideal of the one-dimensional domain $D$, and hence $M$ is maximal.
(2) Consider the localization $D_{M}$ and the $D_{M}$-algebra $A_{M}$. The ring $A_{M}$ cannot be simple, since then $dA_{M}=A_{M}$ for each $d\in D\setminus\{0\}$. This implies that $K=A_{M}\cap K$, and consequently $D\subsetneqq A\cap K$. So, $A_{M}$ contains a nonzero maximal ideal $\mathfrak{M}$. The ideal $\mathfrak{M}\cap D_{M}$ is maximal in $D_{M}$, so we have $MD_{M}=\mathfrak{M}\cap D_{M}$. It follows that the contraction of $\mathfrak{M}$ to $A$ is a proper ideal of $A$ that contains $M$. ∎
@@ 4.1 Building a Prüfer domain inside $K[X]$
Let $D$ be an almost Dedekind domain with finite residue fields. As mentioned in Notation 1.1, $D$ contains a principal ideal domain $D_{0}$ and we use $\mathfrak{p}$ to denote a nonzero prime element of $D_{0}$.
For each nonzero prime $P$ of $D$, let $I_{P}$ be an index set and let $\{V_{P,i}\mid i\in I_{P}\}$ be a collection of valuation overrings of $D[X]$ such that for each $i\in I_{P}$, the valuation $v_{P,i}$ of $K(X)$ extends $v_{P}$, and $V_{P,i}$ has maximal ideal $M_{P,i}$. Let $\mathfrak{p}$ be the prime of $D_{0}$ below $P$. Assume that the following two conditions hold:
@@
    there exists $e_{\mathfrak{p}}\geq 1$ such that for all $P|\mathfrak{p}D_{0}$ and all $i\in I_{P}$, $M_{P,i}^{e_{\mathfrak{p}}}\subseteq\mathfrak{p}V_{P,i}$, and
@@
    there exists $f_{\mathfrak{p}}\geq 1$ such that for all $P|\mathfrak{p}D_{0}$ and all $i\in I_{P}$, $(V_{P,i}/M_{P,i}):(D_{0}/\mathfrak{p}D_{0})]\leq f_{\mathfrak{p}}$.
@@
Since $D$ is assumed to be an almost Dedekind domain, each valuation $v_{P}$ is discrete. The existence of $e_{\mathfrak{p}}$ implies that for each $P|\mathfrak{p}D_{0}$ and each $i\in I_{P}$, there exists a positive integer $\lambda_{P,i}\leq e_{\mathfrak{p}}$ such that $M_{P,i}^{\lambda_{P,i}}=\mathfrak{p}V_{P,i}$. Indeed, the containments $M_{P,i}^{e_{\mathfrak{p}}}\subseteq\mathfrak{p}V_{P,i}\subseteq M_{P,i}$ imply that $\mathfrak{p}V_{P,i}$ is $M_{P,i}$-primary, so by [12, Theorem 17.3(b)] $\mathfrak{p}V_{P,i}$ is equal to a power $M_{P,i}^{\lambda_{P,i}}$ of the maximal ideal $M_{P,i}$ and necessarily $\lambda_{P,i}\leq e_{\mathfrak{p}}$.
@@
Working with the unitary maximal ideals of $\mathscr{D}$ is more complicated. When $\mathfrak{M}$ is unitary, its contraction to $D$ is a nonzero prime $P\subseteq D$, and $P$ itself lies above a prime element $\mathfrak{p}$ of $D_{0}$.
@@
Recall the definitions of the positive integers $e_{\mathfrak{p}}$ and $f_{\mathfrak{p}}$ from Construction 4.1. For each $\mathfrak{p}$, let $s_{\mathfrak{p}}=e_{\mathfrak{p}}!$ and $r_{\mathfrak{p}}=|D_{0}/\mathfrak{p}D_{0}|^{f_{\mathfrak{p}}!}$. Note that $r_{\mathfrak{p}}\geq 2$.
@@
    For each $\mathfrak{p}$ and each $f\in\mathscr{D}$, $\dfrac{(f^{r_{\mathfrak{p}}}-f)^{s_{\mathfrak{p}}}}{\mathfrak{p}}\in\mathscr{D}$.
@@
    Let $\mathfrak{M}$ be a unitary prime ideal of $\mathscr{D}$ above $\mathfrak{p}$. Then, $f^{r_{\mathfrak{p}}}-f\in\mathfrak{M}$ for all $f\in\mathscr{D}$.
@@
For readability, let $r=r_{\mathfrak{p}}$ and $s=s_{\mathfrak{p}}$.
(1) First, we show that $(f^{r}-f)^{s}\in\mathfrak{p}V_{P,i}$ for each $P|\mathfrak{p}D_{0}$ and each $i\in I_{P}$. By construction, $(V_{P,i}/M_{P,i}):(D_{0}/\mathfrak{p}D_{0})]$ divides $f_{\mathfrak{p}}!$ and $e_{\mathfrak{p}}\leq s$. So, $f^{r}-f\in M_{P,i}$, and then $(f^{r}-f)^{s}\in M_{P,i}^{s}\subseteq\mathfrak{p}V_{P,i}$. Thus, $(f^{r}-f)^{s}/\mathfrak{p}\in V_{P,i}$.
Next, if $P^{\prime}$ is a nonzero prime of $D$ that is not above $\mathfrak{p}$, then $\mathfrak{p}$ is invertible in each valuation overring $V^{\prime}$ of $D$ associated to $P^{\prime}$. In this case, $(f^{r}-f)^{s}/\mathfrak{p}\in V^{\prime}$. We conclude that $(f^{r}-f)^{s}/\mathfrak{p}\in\mathscr{D}$.
(2) By part (1), $(f^{r}-f)^{s}\in\mathfrak{p}\mathscr{D}\subseteq\mathfrak{M}$ for each $f\in\mathscr{D}$. Consequently, $f^{r}-f\in\mathfrak{M}$. ∎
@@
Let $f\in\mathscr{D}$. For every $\mathfrak{p}$, every $P|\mathfrak{p}D_{0}$, and each $i\in I_{P}$, one of the following two conditions holds. Either
@@
    $v_{P,i}\Big{(}\dfrac{f^{s_{\mathfrak{p}}}}{\mathfrak{p}^{t}}\Big{)}=0$ for some $t\geq 0$, or
@@
Fix $\mathfrak{p}$, $P$, and $i$. Let $V=V_{P,i}$, $M=M_{P,i}$, $v=v_{P,i}$, and $s=s_{\mathfrak{p}}$. Since $f\in\mathscr{D}$, $v(f)\geq 0$. If $v(f)=0$, then condition (i) holds with $t=0$. So, assume that $v(f)>0$. This means that $f\in M$. If condition (ii) holds, then we are done. So, assume not. Then, there exists some $n\geq 1$ such that $\mathfrak{p}^{n}V\subseteq f^{s}V$. By the same argument of Remark 4.2, $fV=M^{a}$, for some $a\geq 1$, so $f^{s}V=M^{sa}$. Recall from Remark 4.2 that $\mathfrak{p}V=M^{\lambda}$ for some integer $\lambda\geq 1$. Note that $\lambda$ divides $s$ because $\lambda\leq e_{\mathfrak{p}}$. Hence, there exists some $t\geq 1$ such that $sa=\lambda t$, so that $f^{s}V=M^{\lambda t}=\mathfrak{p}^{t}V$. Hence, $v(f^{s})=v(\mathfrak{p}^{t})$, which is (i). ∎
@@
Let $f\in\mathscr{D}$. Assume that for some $\mathfrak{p}$, every $P|\mathfrak{p}D_{0}$, and every $i\in I_{P}$, $f$ satisfies one of the following two conditions. Either
@@
    $v_{P,i}\Big{(}\dfrac{f}{\mathfrak{p}^{t}}\Big{)}>0$ for all $t\geq 0$.
Let $g=\dfrac{f(f^{r_{\mathfrak{p}}-1}-1)^{s_{\mathfrak{p}}}}{\mathfrak{p}}\in K[X]$. Then, $g$ also satisfies (i) or (ii), and $g\in\mathscr{D}$.

### v2

# accepted for publication in the Bulletin of the London Mathematical Society (2026)

@@ 1 Introduction
Integer-valued polynomials have a long history in commutative ring theory. One problem that attracted significant interest [4, 13, 5, 17] was to classify the domains $D$ such that $\textnormal{Int}(D)$ is a Prüfer domain.
Necessary and sufficient conditions on $D$ for this to occur were determined in [18]. Since the late 2000s, attention has turned to rings of the form $\textnormal{Int}_{K}(A)$ and $\textnormal{Int}_{K}(S,A)$ [33, 19, 7, 9, 10, 22, 8, 11, 14, 20, 21]. Here, as in the traditional case where $A=D$, researchers have studied situations in which these rings are or are not Prüfer. For instance, when $D=\mathbb{Z}$ and $A$ is a ring of algebraic integers, it is known that $\textnormal{Int}_{\mathbb{Q}}(A)$ is Prüfer [19, Theorem 3.7]. By contrast, if $D$ is a Dedekind domain with finite residue fields and $A$ is the matrix algebra $A=M_{n}(D)$ with $n\geq 2$, then $\textnormal{Int}_{K}(A)$ is never integrally closed [28, Corollary 3.4], hence cannot be Prüfer.
@@
  • The domain $D$ contains a principal ideal domain $D_{0}$, the structure of which depends on the characteristic of $K$. If $\text{char}(K)=0$, then $D_{0}\cong\mathbb{Z}$. If $\text{char}(K)$ is positive, then $D_{0}\cong F[t]$, the univariate polynomial ring over the finite field $F$ with $\text{char}(K)$ elements.
  • Let $p$ denote a nonzero prime element of $D_{0}$. Explicitly, if $D_{0}\cong\mathbb{Z}$, then $p$ is an integral prime; otherwise, $p$ is a polynomial irreducible over $F$.
@@
  • For each $p$ and each $P$ over $p$, define
    $\quad e_{P,p}=v_{P}(p)\quad\text{ and }\quad f_{P,p}=[(D/P):(D_{0}/pD_{0})].$
  • For each $p$, define
    $E_{p}:=\{e_{P,p}\mid P|pD_{0}\}\quad\text{ and }\quad F_{p}:=\{f_{P,p}\mid P|pD_{0}\}.$
  • We say that $D$ satisfies the double-boundedness condition if, for each $p$, both $E_{p}$ and $F_{p}$ are bounded.
@@ 3 Necessary conditions
In this section, we will establish necessary conditions for $\textnormal{Int}_{K}(A)$ to be Prüfer. Since one of our standard assumptions on $A$ is that $A\cap K=D$, we have $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}(D)$. Thus, if $\textnormal{Int}_{K}(A)$ is Prüfer, then so is $\textnormal{Int}(D)$. If the reader wishes, for the remainder of this section they may consider $D$ to be an ADDB-domain (i.e. $D$ is such that $\textnormal{Int}(D)$ is Prüfer), since this is the assumption we will make in our major theorems. However, we state our intermediary results using only the properties of the integrally closed domain $D$ that are needed to reach the desired conclusion. In fact, a number of these results (Lemmas 3.2, 3.5, 3.6, 3.9; Propositions 3.10, 3.11; Theorem 3.7) hold even when $D$ is not integrally closed.
@@
For any ring $R$, let $\mathscr{J}(R)$ be the Jacobson radical of $R$ and let $\textnormal{Max}(R)$ be the set of maximal ideals of $R$. We recall that $R$ is semiprimitive if $\mathscr{J}(R)=\{0\}$. If $D$ is semiprimitive (which occurs, for instance, when $D=\mathbb{Z}$), we can impose even stronger necessary conditions on $A$ in order for $\textnormal{Int}_{K}(A)$ to be Prüfer. To do this, we first show that in some cases the existence of a nonzero nilpotent element in a simple residue ring of $A$ is enough to force $\textnormal{Int}_{K}(A)$ to not be Prüfer.
@@
Lemma 3.6.
Let $\mathfrak{m}\in\textnormal{Max}(A)$ and let $P=\mathfrak{m}\cap D$.
  1. (1)
    $P$ is a nonzero prime ideal of $D$.
  2. (2)
    If $D$ is one-dimensional, then $P\in\textnormal{Max}(D)$.
  3. (3)
    If $\textnormal{Int}_{K}(A)$ is Prüfer, then $\in\textnormal{Max}(D), PP$ has finite index in $D$, and $\mathfrak{m}$ has finite index in $A$.
Proof.
(1) Certainly, $P$ is a proper ideal of $D$ because $1\notin P$. Let $cd\in P$. Then, $cd\in\mathfrak{m}$, and since both $c$ and $d$ are central in $A$, we have $cAd\subseteq\mathfrak{m}$. The ring $A$ may be noncommutative, but maximal ideals in noncommutative rings are still prime [16, p. 165], so the condition $cAd\subseteq\mathfrak{m}$ implies that either $c\in\mathfrak{m}$ or $d\in\mathfrak{m}$ [16, Prop. 10.2]. Thus, either $c$ or $d$ is in $P$.
Suppose that $P=\{0\}$ and let $R=A/\mathfrak{m}$. Then, $D$ would be isomorphic to a subring of $R$, because
$D=D/(\mathfrak{m}\cap D)\cong(D+\mathfrak{m})/\mathfrak{m}\subseteq R.$
Now, $R$ is a simple ring and $D$ is central in $R$, so for each nonzero $d\in D$ we have $dR=R$. So, $d^{-1}$ exists in $R$. Since $A$ is integral over $D$, this implies that $d^{-1}$ is integral over $D$. It follows that $K$ is integral over $D$. By [1, Proposition 5.7], $D$ is a field. This is a contradiction. Thus, $P\neq\{0\}$.
(2) This follows from (1).
(3) Assume that $\textnormal{Int}_{K}(A)$ is Prüfer. Then, $\textnormal{Int}(D)$ is also Prüfer. By [4, Prop. 6.3], $D$ must be an almost Dedekind domain with finite residue fields. By (2), $P\in\textnormal{Max}(D)$, and hence $D/P$ is finite. Since $A$ is finitely generated as a $D$-module, $A/\mathfrak{m}$ is also finite. ∎
@@
Assume that $A$ contains a maximal ideal $\mathfrak{m}$ such that $A/\mathfrak{m}\cong M_{n}(F)$, where $n\geq 2$ and $F$ is a finite field. Then, $\textnormal{Int}_{K}(A)$ is not Prüfer.
@@
Let $P=D\cap\mathfrak{m}$. By Lemma 3.6(1), $P$ is a nonzero prime ideal of $D$. Suppose that $\textnormal{Int}_{K}(A)$ is Prüfer. Then, as in Lemma 3.6, $D$ is an almost Dedekind domain and $P$ has finite index in $D$. So, $PD_{P}$ is principal. Let $\pi\in D_{P}$ be a generator of $PD_{P}$. Then, $A_{P}/\pi A_{P}$ has $M_{n}(F)$ as a residue ring and $D_{P}:\pi D_{P}]=[D:P]$ is finite. By Lemma 3.5(1), there exists $\alpha\in A_{P}$ such that $\alpha^{2}\in\pi^{2}A_{P}$, but $\alpha\notin\pi A_{P}$. Let $\alpha=a/d$, where $a\in A$ and $d\in D\setminus P$. Then, the ring $\textnormal{Int}_{K}(\{a\},A_{P})$ is not integrally closed, because it contains $X^{2}/(d\pi)^{2}$ but not $X/(d\pi)$. So, $\textnormal{Int}_{K}(\{a\},A_{P})$ is not Prüfer, which is a contradiction because $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}_{K}(\{a\},A_{P})$. ∎
@@
Apply Theorem 3.7 and Lemma 3.6. ∎
@@
Assume that $D$ is one-dimensional. Let $M\in\textnormal{Max}(D)$. Then, there exists a maximal ideal of $A$ containing $MA$.
@@
Consider the localization $D_{M}$ and the $D_{M}$-algebra $A_{M}$. The ring $A_{M}$ cannot be simple, since then $dA_{M}=A_{M}$ for each $d\in D\setminus\{0\}$. This implies that $K=A_{M}\cap K$, and consequently $D\subsetneqq A\cap K$. So, $A_{M}$ contains a nonzero maximal ideal $\mathfrak{M}$. The ideal $\mathfrak{M}\cap D_{M}$ is maximal in $D_{M}$ by Lemma 3.6(2), so we have $MD_{M}=\mathfrak{M}\cap D_{M}$. It follows that the contraction of $\mathfrak{M}$ to $A$ is a proper ideal of $A$ that contains $M$. ∎
@@ 4.1 Building a Prüfer domain inside $K[X]$
Let $D$ be an almost Dedekind domain with finite residue fields. As mentioned in Notation 1.1, $D$ contains a principal ideal domain $D_{0}$ and we use $p$ to denote a nonzero prime element of $D_{0}$.
For each nonzero prime $P$ of $D$, let $I_{P}$ be an index set and let $\{V_{P,i}\mid i\in I_{P}\}$ be a collection of valuation overrings of $D[X]$ such that for each $i\in I_{P}$, the valuation $v_{P,i}$ of $K(X)$ extends $v_{P}$, and $V_{P,i}$ has maximal ideal $M_{P,i}$. Let $p$ be the prime of $D_{0}$ below $P$. Assume that the following two conditions hold:
@@
    there exists $e_{p}\geq 1$ such that for all $P|pD_{0}$ and all $i\in I_{P}$, $M_{P,i}^{e_{p}}\subseteq pV_{P,i}$, and
@@
    there exists $f_{p}\geq 1$ such that for all $P|pD_{0}$ and all $i\in I_{P}$, $(V_{P,i}/M_{P,i}):(D_{0}/pD_{0})]\leq f_{p}$.
@@
Since $D$ is assumed to be an almost Dedekind domain, each valuation $v_{P}$ is discrete. The existence of $e_{p}$ implies that for each $P|pD_{0}$ and each $i\in I_{P}$, there exists a positive integer $\lambda_{P,i}\leq e_{p}$ such that $M_{P,i}^{\lambda_{P,i}}=pV_{P,i}$. Indeed, the containments $M_{P,i}^{e_{p}}\subseteq pV_{P,i}\subseteq M_{P,i}$ imply that $pV_{P,i}$ is $M_{P,i}$-primary, so by [12, Theorem 17.3(b)] $pV_{P,i}$ is equal to a power $M_{P,i}^{\lambda_{P,i}}$ of the maximal ideal $M_{P,i}$ and necessarily $\lambda_{P,i}\leq e_{p}$.
@@
Working with the unitary maximal ideals of $\mathscr{D}$ is more complicated. When $\mathfrak{M}$ is unitary, its contraction to $D$ is a nonzero prime $P\subseteq D$, and $P$ itself lies above a prime element $p$ of $D_{0}$.
@@
Recall the definitions of the positive integers $e_{p}$ and $f_{p}$ from Construction 4.1. For each $p$, let $s_{p}=e_{p}!$ and $r_{p}=|D_{0}/pD_{0}|^{f_{p}!}$. Note that $r_{p}\geq 2$.
@@
    For each $p$ and each $f\in\mathscr{D}$, $\dfrac{(f^{r_{p}}-f)^{s_{p}}}{p}\in\mathscr{D}$.
@@
    Let $\mathfrak{M}$ be a unitary prime ideal of $\mathscr{D}$ above $p$. Then, $f^{r_{p}}-f\in\mathfrak{M}$ for all $f\in\mathscr{D}$.
@@
For readability, let $r=r_{p}$ and $s=s_{p}$.
(1) First, we show that $(f^{r}-f)^{s}\in pV_{P,i}$ for each $P|pD_{0}$ and each $i\in I_{P}$. By construction, $(V_{P,i}/M_{P,i}):(D_{0}/pD_{0})]$ divides $f_{p}!$ and $e_{p}\leq s$. So, $f^{r}-f\in M_{P,i}$, and then $(f^{r}-f)^{s}\in M_{P,i}^{s}\subseteq pV_{P,i}$. Thus, $(f^{r}-f)^{s}/p\in V_{P,i}$.
Next, if $P^{\prime}$ is a nonzero prime of $D$ that is not above $p$, then $p$ is invertible in each valuation overring $V^{\prime}$ of $D$ associated to $P^{\prime}$. In this case, $(f^{r}-f)^{s}/p\in V^{\prime}$. We conclude that $(f^{r}-f)^{s}/p\in\mathscr{D}$.
(2) By part (1), $(f^{r}-f)^{s}\in p\mathscr{D}\subseteq\mathfrak{M}$ for each $f\in\mathscr{D}$. Consequently, $f^{r}-f\in\mathfrak{M}$. ∎
@@
Let $f\in\mathscr{D}$. For every $p$, every $P|pD_{0}$, and each $i\in I_{P}$, one of the following two conditions holds. Either
@@
    $v_{P,i}\Big(\dfrac{f^{s_{p}}}{p^{t}}\Big)=0$ for some $t\geq 0$, or
@@
Fix $p$, $P$, and $i$. Let $V=V_{P,i}$, $M=M_{P,i}$, $v=v_{P,i}$, and $s=s_{p}$. Since $f\in\mathscr{D}$, $v(f)\geq 0$. If $v(f)=0$, then condition (i) holds with $t=0$. So, assume that $v(f)>0$. This means that $f\in M$. If condition (ii) holds, then we are done. So, assume not. Then, there exists some $n\geq 1$ such that $p^{n}V\subseteq f^{s}V$. By the same argument of Remark 4.2, $fV=M^{a}$, for some $a\geq 1$, so $f^{s}V=M^{sa}$. Recall from Remark 4.2 that $pV=M^{\lambda}$ for some integer $\lambda\geq 1$. Note that $\lambda$ divides $s$ because $\lambda\leq e_{p}$. Hence, there exists some $t\geq 1$ such that $sa=\lambda t$, so that $f^{s}V=M^{\lambda t}=p^{t}V$. Hence, $v(f^{s})=v(p^{t})$, which is (i). ∎
@@
Let $f\in\mathscr{D}$. Assume that for some $p$, every $P|pD_{0}$, and every $i\in I_{P}$, $f$ satisfies one of the following two conditions. Either
@@
    $v_{P,i}\Big(\dfrac{f}{p^{t}}\Big)>0$ for all $t\geq 0$.
Let $g=\dfrac{f(f^{r_{p}-1}-1)^{s_{p}}}{p}\in K[X]$. Then, $g$ also satisfies (i) or (ii), and $g\in\mathscr{D}$.